Two people carry a heavy object by placing it on a uniform board that is 2 meters long and weighs 200N. One person lifts at one end with a force of 400N, and the other lifts the opposites end with a force of 600N. Find the weight and the location of the heavy object. 


1
Expert's answer
2021-07-01T04:57:56-0400

Solution.

"l=2m;"

"F_1=400N;"

"F_2=600N;"

"F_3=200N;"

"F_1+F_2=F+F_3\implies F=F_1+F_2-F_3;"

"F=400N+600N-200N=800N;"

"T_{net}=Fx+F_3(l\/2)-F_2l=0;"

"800x+200-1200=0;"

"800x=1000;"

"x=1000\/800=1.25m;"

Answer: "F=800N;"

"x=1.25m."